The Gillman Group issued $900,000 of 13% bonds on June 30 Year 1, for $967,707. The bonds were dated on June 30 and mature on June 30, Year 21 (20 years). The market rate for bonds of similar risk and maturity is 12%. Interest is payable semiannually on June 30 and December 31.

Instructions:

A. Prepare the journal entry to record the issuance of the bonds on June 30, Year 1.

B. Prepare the journal entry to record interest expense on December 31, Year 1 on the effective-

     interest method.  

C. Prepare the journal entry to record interest expense on June 30, Year 2 on the effective-

     interest method.
Answer #1

Date June 30 Year 1 Issue Entry 1. Record the Bond Issue General Journal Cash Premium on Bonds Payable Bonds Payable Record t900000 967707 58500 58062 58036 Face Value Isssue Price Cash interest =900000*13%*6/12 Interest Expenses=967707*12%*6/12 9672